OpenStudy (anonymous):

Find all solutions to the equation in radians: 2(sqrt3) - 4 = -4 - 2tan(theta/3)

OpenStudy (michele_laino):

I can rewrite your equation as below: \[2\sqrt{3}=-2 \tan(\theta/3)\] so: \[\tan(\theta/3)=-\sqrt{3}\] whose solution are:  \[\frac{ \theta }{ 3 }=\frac{ 5 \pi }{ 3 }+k*\pi\] so: \[\theta=5 \pi+3*k*\pi\] where: \[k=0, \pm1, \pm2, \pm3....\]

OpenStudy (anonymous):

I'm confused... where did the 5pi/3 come from? The unit circle?

OpenStudy (michele_laino):

yes!  you can choose: \[\frac{ \theta }{ 3 }=\frac{ 2 \pi }{ 3 }+k*\pi\] or \[\frac{ \theta }{ 3 }=\frac{ 5 \pi }{ 3 }+k*\pi\] as I did!